# DSwinIR: Rethinking Window-based Attention for Image Restoration

### v2

@@ Abstract
Image restoration has witnessed significant advancements with the development of deep learning models. Especially Transformer-based models, particularly those leveraging window-based self-attention, have become a dominant force in image restoration. However, their performance is fundamentally constrained by the rigid, non-overlapping window partitioning scheme, which leads to two critical limitations: insufficient feature interaction across window boundaries and content-agnostic receptive fields that cannot adapt to diverse image structures. Existing methods often rely on heuristic patterns to mitigate these issues, rather than addressing the root cause. In this paper, we propose the Deformable Sliding Window Transformer (DSwinIR), a new foundational backbone architecture that systematically overcomes these limitations. At the heart of DSwinIR is the proposed novel Deformable Sliding Window (DSwin) Attention. This mechanism introduces two fundamental innovations. First, it replaces the rigid partitioning with a token-centric sliding window paradigm, ensuring seamless cross-window information flow and effectively eliminating boundary artifacts. Second, it incorporates a content-aware deformable sampling strategy, which allows the attention mechanism to learn data-dependent offsets and dynamically shape its receptive fields to focus on the most informative image regions. This synthesis endows the model with both strong locality-aware inductive biases and powerful, adaptive long-range modeling capabilities. Extensive experiments show that DSwinIR sets a new state-of-the-art across a wide spectrum of image restoration tasks. For instance, in all-in-one restoration, our DSwinIR surpasses the most recent backbone GridFormer by over 0.53 dB on the three-task benchmark and a remarkable 0.86 dB on the five-task benchmark. The code and pre-trained models are available at https://github.com/Aitical/DSwinIR.
Index Terms:
All-in-One Image Restoration, Image Deraining, Image Dehazing, Image Denoising, Vision Transformer, Deformable Attention
Figure 1: Comparative analysis of feature extraction mechanisms with an anchor token (marked by $\star$) as the reference point. (a) Vanilla convolution applies a fixed sampling pattern, leveraging neighborhood features. (b) Deformable convolution introduces adaptive sampling locations based on content, enabling more effective feature integration from relevant regions. (c) Window attention suffers from boundary constraints where anchor tokens near window edges (especially corners) have limited receptive fields, resulting in suboptimal feature extraction. (d) Our proposed Deformable Sliding Window (DSwin) extends window attention with token-centric paradigm and the content-adaptive reception field, ensuring robust feature aggregation for anchor tokens.
@@ I Introduction
Image restoration, a fundamental and long-standing challenge in computer vision, aims to recover pristine, high-quality images from observations degraded by various factors such as noise, haze, rain, or blur [1, 2, 3, 4]. In the last decades, the advent of deep learning, particularly Convolutional Neural Networks (CNNs), has heralded a new era in this field, with architectures achieving remarkable performance on specific restoration tasks [5, 6]. More recently, Vision Transformers (ViTs) [7] have emerged as a powerful alternative, showcasing exceptional capabilities in modeling long-range dependencies, a crucial aspect for handling complex and spatially extensive degradations.
@@
The practical application of vanilla Transformers to image restoration, however, is hampered by the quadratic computational complexity of the global self-attention mechanism with respect to input resolution. To surmount this obstacle, window-based self-attention, popularized by the Swin Transformer [8], has become the de facto standard for restoration models like SwinIR [9], and Uformer [10]. By restricting attention computation to small local windows, these models achieve a favorable balance between performance and efficiency. Yet, this very design choice—the rigid, non-overlapping window partition—is also the architecture’s Achilles’ heel. It introduces two fundamental, intertwined limitations: 1) Boundary Information Truncation: Tokens near window edges, especially at the corners, have their contextual information severely truncated, leading to suboptimal and spatially-variant feature extraction. 2) Inflexible Receptive Fields: The fixed, square-shaped attention pattern is applied uniformly across the image, failing to adapt to the diverse and anisotropic nature of image content and degradation patterns. Several subsequent works have attempted to alleviate these issues. The original Swin Transformer introduced a shifted-window mechanism. Others have proposed more sophisticated, fixed interaction patterns, such as cross-shaped windows [11], axial stripes [12], or sparse token sampling [13]. While these methods have demonstrably improved performance by enhancing inter-window communication, they are essentially heuristic patches. They attempt to compensate for one fixed pattern (non-overlapping windows) by introducing other, more complex fixed patterns. They fall short of addressing the core, underlying principle: the shape and scope of an attention field should be dynamically determined by the image content itself, not by a predefined, static template.
In this work, we argue for a principled departure from such heuristic designs. We posit that a truly effective attention mechanism for image restoration must be both token-centric and content-aware. Inspired by this, we introduce a novel attention mechanism, Deformable Sliding Window (DSwin) Attention, which fundamentally redesigns window-based attention, as illustrated in Fig. 1. First, to address boundary truncation and foster seamless information flow, we embrace a token-centric sliding window paradigm. Analogous to the sliding operation of a convolutional kernel, this ensures that every token is at the center of a complete contextual window, naturally creating overlaps and robustly integrating neighborhood features. Second, to tackle the inflexible receptive field, we introduce a content-aware deformable mechanism. Drawing inspiration from deformable convolutions [14], our mechanism learns to predict spatial offsets based on local features, allowing the attention window to dynamically warp and focus on the most salient regions, thus creating effective, data-dependent receptive fields. Building upon this core innovation, we present the Deformable Sliding Window Transformer for Image Restoration (DSwinIR). This new backbone architecture systematically integrates the DSwin attention module, offering a more powerful and flexible foundation for image restoration.
In recent years, the development of all-in-one image restoration models—capable of handling multiple types of degradation within a unified framework—has gained increasing attention. Such models have largely followed two distinct philosophies: one line of work focuses on using task-specific prompts to guide a generic backbone [15, 16], while the other emphasizes evolving the backbone architecture itself to achieve broader restoration capabilities. However, despite the growing interest in prompt-based approaches, we argue that strengthening the backbone architecture remains a fundamental and underexplored avenue. Our work represents a significant advancement for the latter philosophy, demonstrating that a sufficiently powerful and adaptive backbone alone can achieve state-of-the-art performance across diverse tasks, without relying on external prompt-based guidance. This focus on architectural fundamentals is crucial for ensuring both generalization and long-term progress in the field.
We conduct extensive evaluations across diverse image restoration tasks, spanning both all-in-one multiple degradation scenarios and specialized single-task settings. As demonstrated in Fig.2, DSwinIR delivers substantial improvements of 2.1 dB and 1.3 dB in synthetic and real-world deweathering tasks, respectively. Moreover, our approach establishes new state-of-the-art performance on three-task and five-task degradation benchmarks, outperforming previous methods by approximately 0.7 dB and 0.9 dB. For single-task restoration, DSwinIR surpasses the current leading method DRSformer[13] by 0.62 dB on the challenging real-world deraining SPA dataset [17].
@@
  • We propose a novel Deformable Sliding Window (DSwin) Attention mechanism. It pioneers a token-centric, content-adaptive paradigm for window-based attention, fundamentally solving the boundary truncation and inflexible receptive field limitations of prior art.
  • We develop DSwinIR, a new and powerful backbone architecture for image restoration built upon the DSwin attention module. The architecture cohesively integrates multi-scale feature extraction in both the attention and feed-forward network modules, further enhancing its representation capacity.
  • Through extensive experiments on a diverse array of benchmarks, including multi-degradation all-in-one settings and specialized single-task restoration, we demonstrate that DSwinIR consistently and significantly outperforms existing state-of-the-art methods, affirming the superiority of our architectural innovations.
The remainder of this paper is organized as follows. Section II provides a comprehensive review of the image restoration literature, charting the evolution from specialized single-task models to modern all-in-one frameworks. Section III elaborates on our proposed method, detailing the principled progression from the limitations of standard window attention to our novel Deformable Sliding Window (DSwin) attention, and its integration into the complete DSwinIR architecture. Section IV presents extensive experimental validation on a wide array of benchmarks, including quantitative comparisons, qualitative results, in-depth ablation studies. Finally, Section V concludes the paper by summarizing our key contributions and their broader significance for the field.
@@ II Related Work
Image restoration has evolved from specialized, single-task models to highly versatile all-in-one frameworks. In this section, we chart this evolution to contextualize our contribution. We begin by reviewing the progress in single-task restoration, covering key advancements from the CNN era to the rise of Transformers and other emerging architectures. We then provide a comprehensive analysis of the all-in-one restoration landscape, examining the various innovative strategies developed to tackle multi-degradation scenarios. This analysis will situate our work within the broader field and highlight the unique perspective we bring to architectural design.
@@ II-A Single-Task Image Restoration
Deep learning, particularly with the rise of Convolutional Neural Networks (CNNs), has long been the cornerstone of image restoration. Leveraging a data-driven paradigm, CNNs have proven highly effective compared to classical approaches [18], establishing strong foundations in tasks such as image super-resolution [19], image denoising [5], image deraining [20], and image dehazing [21]. Subsequently, a multitude of more sophisticated studies emerged, for example, models incorporating residual or dense connections [22, 23], as well as those exploring channel attention [24] and non-local attention mechanisms [25]. Within specific domains, landmark contributions have guided the field; for instance, in image deraining, Jiang et al. proposed MSPFN [26], which introduced a progressive learning paradigm that inspired a series of subsequent works [27, 28, 29]. Similarly, image dehazing has also seen many representative studies, such as FFANet [30], and more recent explorations with novel network designs [31, 32]. In parallel, researchers have also focused on developing general restoration architectures applicable to multiple tasks, including MPRNet [33], NAFNet [34], FSNet [35], IRNeXT [36], ConvIR [37], and MB-TaylorFormer [38]. However, these powerful models still typically require separate training for each individual task.
@@
Concurrently, to better model the complex, non-local dependencies often present in degradations, Vision Transformers (ViTs) [7] have garnered significant attention for their powerful dynamic modeling capabilities. While early models like IPT [39] were computationally intensive, the window-based local attention of the Swin Transformer [8] provided an efficient and effective foundation. This spawned a new generation of influential backbones, including SwinIR [9] and Uformer [10]. These advancements were also channeled into powerful task-specific transformers like IDT [40] for image deraining and Dehazeformer [41] for image dehazing. A significant thread of subsequent research has focused extensively on refining the window-based mechanism. For example, HAT [42] enlarged window sizes and used overlaps to enhance cross-window communication, while ELAN [43] and SRFormer [44] addressed computational redundancy. More sophisticated attention patterns were also developed, such as the sparse attention in DRSFormer [13], the strip attention in StripFormer [12], the transposed self-attention in Restormer [45], the integration of an adaptive token dictionary [46], and progressive focused attention [47]. Furthermore, hybrid models like GRL [48] and X-Restormer [49] have explored the fruitful combination of convolutional and attention-based feature extraction.
@@
Alongside these core architectural advancements, several complementary research lines have gained traction. Emerging paradigms like state-space models, exemplified by MambaIR [50, 51], offer a compelling alternative with linear complexity, and Wang et al. [52] exploited the vision autoregressive model for all-in-one image restoration . Moreover, researchers are exploring other critical facets of image restoration, such as estimating model uncertainty [53, 54, 55, 56] and leveraging self-supervised pre-training with contrastive learning to learn robust image priors without paired data [57, 58, 59, 60]. These diverse efforts collectively enrich the field and provide a strong foundation for tackling more complex restoration scenarios.
II-B All-in-One Restoration: The Quest for a Unified Model
@@
Addressing the challenges of all-in-one restoration, where a single model must gracefully handle diverse degradations, has inspired a variety of innovative strategies. A significant body of work has focused on developing sophisticated prompting mechanisms to guide general-purpose backbones, adapting their behavior to specific tasks. This approach often involves learning adaptive prompts directly from the data. Influential works like AirNet [61] and PromptIR [15] demonstrated how a learned degradation representation can effectively modulate a restoration network. To enhance the discriminability of these prompts, subsequent methods introduced explicit regularization, such as the orthogonality constraints in IDR [62] and the frequency-domain regularization in AdaIR [63]. Others have explored more expressive prompt representations, for instance, by modeling the prompt as a learned distribution as in DPPD [64]. A parallel thread in prompt-based learning injects richer semantic knowledge from large, pre-trained multi-modal models. Methods like DA-CLIP [65], TextualDegRemoval [66], Perceive-IR [67], and OnmiFusion [68] leverage models like CLIP [69] to provide textual or multi-modal guidance. More recently, architectures leveraging Mixture-of-Experts (MoE), such as MoCE-IR [70], have been exploited to select specialized parameters for different restoration tasks.
@@
Complementing the work on prompt-based guidance, other research investigates the optimization challenges inherent in training a single model on diverse data. From a multi-task learning perspective, methods like [71] have explored techniques to mitigate task conflicts, while others like [72] focus on correcting for data and task bias during joint training. Similarly, contrastive learning has been adapted to the all-in-one setting to help the model better decouple degradation-specific features from content-invariant features [73, 74].

### v3

@@ Abstract
Image restoration has witnessed significant advancements with the development of deep learning models. Transformer-based models, particularly those using window-based self-attention, have become a dominant force. However, their performance is constrained by the rigid, non-overlapping window partitioning scheme, which leads to insufficient feature interaction across windows and limited receptive fields. This highlights the need for more adaptive and flexible attention mechanisms. In this paper, we propose the Deformable Sliding Window Transformer for Image Restoration (DSwinIR), a new attention mechanism: the Deformable Sliding Window (DSwin) Attention. This mechanism introduces a token-centric and content-aware paradigm that moves beyond the grid and fixed window partition. It comprises two complementary components. First, it replaces the rigid partitioning with a token-centric sliding window paradigm, making it effective at eliminating boundary artifacts. Second, it incorporates a content-aware deformable sampling strategy, which allows the attention mechanism to learn data-dependent offsets and actively shape its receptive field to focus on the most informative image regions. Extensive experiments show that DSwinIR achieves strong results, including state‑of‑the‑art performance on several evaluated benchmarks. For instance, in all-in-one image restoration, our DSwinIR surpasses the most recent backbone GridFormer by 0.53 dB on the three-task benchmark and 0.87 dB on the five-task benchmark. The code and pre-trained models are available at https://github.com/Aitical/DSwinIR.
Figure 1: Comparative analysis of feature extraction mechanisms with an anchor token (marked by $\star$) as the reference point. (a) Vanilla convolution applies a fixed sampling pattern, leveraging neighborhood features. (b) Deformable convolution introduces adaptive sampling locations based on content, enabling more effective feature integration from relevant regions. (c) Window attention suffers from boundary constraints where anchor tokens near window edges (especially corners) have limited receptive fields. (d) The proposed Deformable Sliding Window (DSwin) attention extends window attention with a token-centric paradigm and the content-aware sampling, bringing robust feature aggregation for anchor tokens.
@@ I Introduction
Image restoration aims to recover high‑quality images from inputs degraded by noise, rain, haze, blur, low light, and other factors [survey_denoise_Michael2023, survey_dehaze_gui2023, survey_derain_chen2023, Zhang22blurSurvey, survey_all_in_one_jiang2024]. Deep learning has led to substantial progress, initially with CNNs [DnCNN, zamir2021multi] and more recently, with Vision Transformers (ViTs) [ViT_Alexey_2021] that model long‑range dependencies.
@@
The practical use of vanilla Transformers for image restoration is impeded by the quadratic cost of global self-attention with respect to input resolution. Window-based self-attention [SwinT_LiuL00W0LG21], adopted in SwinIR [SwinIR] and Uformer [Uformer], has therefore become the de facto choice, striking a favorable trade-off between fidelity and efficiency by restricting computation to local, non-overlapping windows. It introduces two salient limitations: 1) Boundary Context Truncation. Consider a token at a window corner: its receptive field is clipped precisely where informative context often lies across the boundary. As a result, its representation collapses toward its nearest in-window neighbors rather than the true cross-window structure it belongs to, producing suboptimal, biased features. 2) Inflexible Receptive Fields: The fixed, square-shaped attention pattern is applied uniformly across the image, failing to adapt to the diverse and anisotropic nature of image content and degradation patterns. Subsequent efforts introduce hand-crafted interaction patterns, such as cross-shaped windows [CAT_ChenZGzKY22], axial stripes [StripFormer_TsaiPLTL22], and sparse token sampling [DRSformer]. While effective, these essentially replace one fixed interaction pattern with another, leaving the core limitation unresolved. They fall short of addressing the core, underlying principle: the shape and scope of an attention field should be dynamically determined by the image content itself, not by a predefined, static template.
In this work, we move beyond such predefined, static patterns and advocate that an effective attention mechanism for image restoration is both token-centric and content-aware. Motivated by this view, we introduce Deformable Sliding Window (DSwin) Attention, a redesign of windowed attention (Figure 1). First, to strengthen cross‑window interaction, we embrace a token-centric sliding window paradigm: every token attends within its own centered neighborhood, creating overlaps that reduce boundary effects. Second, to tackle the inflexible receptive field, we introduce a content-aware deformable mechanism: learned offsets adapt the sampling pattern to image structure (e.g., diagonal rain streaks or object contours). Building upon core components, we present the Deformable Sliding Window Transformer for Image Restoration (DSwinIR). It integrates the DSwin attention module, offering a flexible and robust backbone for diverse image restoration tasks.
In recent years, the development of all-in-one image restoration models, handling multiple types of degradation within a unified framework, has gained increasing attention [survey_all_in_one_jiang2024]. We evaluate DSwinIR in the all‑in‑one setting. Strengthening the backbone architecture offers an underexplored alternative to prompt‑based approaches [PromptIR, cui2025adair]. Empirically, DSwinIR matches or exceeds several prompt‑based methods on our benchmarks. We conduct extensive evaluations across diverse image restoration tasks, spanning both all-in-one multiple degradation scenarios and specialized single-task settings. As illustrated in Figure 2, DSwinIR delivers substantial improvements across diverse benchmarks.
@@
  • We propose a novel Deformable Sliding Window (DSwin) Attention mechanism. It adopts a token‑centric, content‑aware design that addresses boundary truncation and inflexible receptive fields in window‑based attention.
  • We develop DSwinIR, a new and general backbone architecture for image restoration built upon the DSwin attention module. The architecture cohesively integrates multi-scale feature extraction in both the attention and feed-forward network modules, further enhancing its representation capacity.
  • Through extensive experiments on a diverse array of benchmarks, including all-in-one and single-task restoration scenarios, we demonstrate that DSwinIR consistently achieves competitive or superior performance compared to existing methods on the evaluated datasets.
The remainder of this paper is organized as follows. Section II provides a comprehensive review of the image restoration literature. Section III elaborates on our proposed method. Section IV presents extensive experimental validation on a wide array of benchmarks. Finally, Section V concludes the paper by summarizing our key contributions and their broader significance for the field.
@@ II Related Work
Image restoration, a classic and long-standing challenge in computer vision, has seen decades of research [Digital_IR_Survey]. It has a wide-ranging impact on various downstream tasks, from enhancing visual data for autonomous driving [Almalioglu2022weatherDrive] and improving the robustness of person re-identification [Xu2022Rankinrank] to enabling one-click image correction for cellular segmentation in bioinformatics [Stringer2025Cellpose3]. Nowadays, the advent of deep learning has dramatically accelerated progress in this field [transformer_IR], leading to remarkable breakthroughs in specialized tasks like denoising[survey_denoise_Michael2023], dehazing [survey_dehaze_gui2023], and deraining[survey_derain_chen2023]. More recently, the field’s focus has evolved from single-task models to highly versatile all-in-one frameworks [survey_all_in_one_jiang2024], which are more practical for real-world scenarios involving complex, mixed degradations. In this section, we chart this evolution to contextualize our contribution.
@@ II-A Single-Task Image Restoration
Deep learning, particularly with the rise of Convolutional Neural Networks (CNNs), has long been the cornerstone of image restoration. Leveraging a data-driven paradigm, CNNs have proven highly effective compared to classical approaches [Digital_IR_Survey], establishing strong foundations in tasks such as image super-resolution [SRCNN], image denoising [DnCNN], image deraining [fu2017deraincnn], and image dehazing [DehazeNet]. Subsequently, numerous more sophisticated studies emerged, as well as those incorporating residual or dense connections [EDSR, RDN], as well as those exploring channel attention [RCAN] and non-local attention mechanisms [NLSN]. Within specific domains, landmark contributions have guided the field; for instance, in image deraining, Jiang et al. proposed MSPFN [jiang2020multi], which introduced a progressive learning paradigm that inspired a series of subsequent works [Ni2021derain_cvpr, Zhou2021continual_derain, chen2024nerfDerain]. Image dehazing has also seen many representative studies, such as FFANet [Qin2020ffanet], and more recent explorations with novel network designs [cui2024FocalNet, li2025pami_dehaze]. Similarly, Feijoo et al. [Feijoo2025DarkIR] exploit a light-weight yet powerful model for low-light image enhancement in the proposed DarkIR. In parallel, researchers have also focused on developing general restoration architectures applicable to multiple tasks, including MPRNet [ZamirA2021Mprnet], NAFNet [chen2022simple], FSNet [CuiRCK24], IRNeXT [cui2023irnext], ConvIR [Cui2024ConvIR], and MB-TaylorFormer [Jin2025mbtaylor]. However, these powerful models still typically require separate training for each individual task.
@@
Concurrently, to better model the complex, non-local dependencies often present in degradations, Vision Transformers (ViTs) [ViT_Alexey_2021] have garnered significant attention for their powerful dynamic modeling capabilities. While early models like IPT [IPT] were computationally intensive, the window-based local attention of the Swin Transformer [SwinT_LiuL00W0LG21] provided an efficient and effective foundation. This spawned a new generation of influential backbones, including SwinIR [SwinIR] and Uformer [Uformer]. These advancements were also channeled into powerful task-specific transformers like IDT [Xiao2023IDT] for image deraining and Dehazeformer [Song2023DehazeFormer] for image dehazing. A significant thread of subsequent research has focused extensively on refining the window-based mechanism. For example, HAT [HAT_ChenWZ0D23] enlarged window sizes and used overlaps to enhance cross-window communication, while ELAN [zhang2022elan] and SRFormer [zhou2023srformer] addressed computational redundancy. More sophisticated attention patterns were also developed, such as the sparse attention in DRSFormer [DRSformer], the strip attention in StripFormer [StripFormer_TsaiPLTL22], the transposed self-attention in Restormer [Restormer], the integration of an adaptive token dictionary [Zhang2024ATD], and progressive focused attention [Long2025progressiveSA]. Furthermore, hybrid models like GRL [Li2023GRL] and X-Restormer [Chen2024xRestormer] have explored the fruitful combination of convolutional and attention-based feature extraction.
@@
Alongside these core architectural advancements, several complementary research lines have gained traction. Emerging paradigms like state-space models, exemplified by MambaIR [MambaIR, guo2025mambav2], offer a practical solution for image restoration tasks. Meanwhile, Shi et al. [Shi2025VmambaIR] introduce the VmambaIR by the proposed omni-scan, and Li et al. [Li2025MaIR] propose the MaIR with the local S-shape scanning and sequential shuffle attention. Besides, Wang et al. [wang2025navigatingimagerestorationvars] exploited the vision autoregressive model for all-in-one image restoration. Moreover, researchers are exploring other critical facets of image restoration, such as estimating model uncertainty [Ning2021uncertaintydrivenSR, Hong2022uncertaintydehazing, Huang2023uncertaintyDerain, Qiao2024uncertaintyDiffHaze] and leveraging self-supervised pre-training with contrastive learning to derive robust image priors without paired data [Wu2021contrastiveHaze, Zheng2023Curriculardehaze, chen2022unpaircontrastivederain, wu2024modelcontrastive]. These diverse efforts collectively enrich the field and provide a strong foundation for tackling more complex restoration scenarios.
II-B All-in-One Image Restoration: The Quest for a Unified Model
@@
Addressing the challenges of all-in-one image restoration, where a single model must gracefully handle diverse degradations, has inspired a variety of innovative strategies [survey_all_in_one_jiang2024]. A significant body of work has focused on developing sophisticated prompting mechanisms to guide general-purpose backbones, adapting their behavior to specific tasks. This approach often involves learning adaptive prompts directly from the data. Influential works like AirNet [AirNet] and PromptIR [PromptIR] demonstrated how a learned degradation representation can effectively modulate a restoration network. To enhance the discriminability of these prompts, subsequent methods introduced explicit regularization, such as the orthogonality constraints in IDR [IDR] and the frequency-domain regularization in AdaIR [cui2025adair]. Rajagopalan et al. [Rajagopalan2025AWRaCLe] exploit the task prompt representation with the in-context learning paradigm. Others have explored more expressive prompt representations, for instance, by modeling the prompt as a learned distribution as in DPPD [wu2025dppd]. A parallel thread in prompt-based learning injects richer semantic knowledge from large, pre-trained multimodal models. Zheng et al. [Zeng2025VLUNet] propose the deep unfolding model with pretrained multimodal prior. Methods like DA-CLIP [Luo2024DA-CLIP], TextualDegRemoval [texture], Perceive-IR [zhang2024perceive], and OmniFuse [Zhang2025onmifusion] leverage models like CLIP [Radford2021CLIP] to provide textual or multimodal guidance. More recently, architectures leveraging Mixture-of-Experts (MoE), such as MoCE-IR [Zamfir2025MOCEIR], have been exploited to select specialized parameters for different restoration tasks.
@@
Complementing the work on prompt-based guidance, another significant research direction investigates generative diffusion models [weather_diffusion, DTPM_0001CCXQL024]. For instance, Özdenizci et al. [weather_diffusion] proposed the first diffusion-based deweathering model. Following this, Ye et al. [DTPM_0001CCXQL024] enhanced the diffusion framework with task-aware adapters to better address multiple degradation tasks. Besides, some research investigates the optimization challenges inherent in training a single model on diverse data. From a multi-task learning perspective, methods like [wu2024harmony] have explored techniques to mitigate task conflicts, while others like [wu2025debiased] focus on correcting for data and task bias during joint training. Similarly, contrastive learning has been adapted to the all-in-one setting to help the model better decouple degradation-specific features from content-invariant features [tang2025contrastive, wu2025cpl].